At a football game, 200 souvenirs were sold, some at $1.50 each and the rest at $2.50 each. If the proceeds from the sales totaled $380, how many of each kind were sold?
Answer by checkley77(12844)

2.50x+1.50(200-x)=380
2.50x+300-1.50x=380
x=380-300
x=80 $2.50 souvenirs were sold.
200-80=120 $1.50 souvenirs were sold.
Proof:
2.50*80+1.50*120=380
200+180=380
380=380